# Numerical validation of scaling laws for stratified turbulence

### v1

@@ 3 The effects of intermittency
To gain insight into the applicability of the predicted scalings, we examine the actual flow field more closely. Figure 3 shows snapshots of $u$ and $w$ in two different high-resolution DNS at $Pe=100$ and $Re=1000$. The left panels show a more weakly stratified case with $Fr\simeq 0.18$ and the right panels show a more strongly stratified case with $Fr\simeq 0.058$. It is clear that while the $Fr\simeq 0.18$ case is fully turbulent, the more strongly stratified case is not, but exhibits instead spatial intermittency in as much as the turbulence is localized to small ‘patches’ (see e.g. the regions of high $|w|$ in the bottom right panel). Similar findings were reported by Cope et al., (2020) in their low $Pe$ simulations in the regime that they named ‘stratified intermittent’ (see their figure 6).
Figure 3: DNS snapshots of $u$ (top row) and $w$ (bottom row) in the $y=0$ plane for $Re=1000$, $Pe=100$. The left column shows a more weakly stratified simulation with $Fr=30^{-1/2}\simeq 0.18$ and the right column shows a more strongly stratified simulation with $Fr=300^{-1/2}\simeq 0.058$. Note the spatially intermittent nature of the turbulence in the latter.
@@
These snapshots demonstrate two very important aspects of strongly stratified turbulence that have direct implications for asymptotic analyses. First, they illustrate the coexistence of large and small horizontal scales, with $u$ dominated by large scales with subdominant small scales, and $w$ dominated by small scales with subdominant large scales (cf. Riley and Lindborg,, 2012). This ordering is a cornerstone of the MSA theories of Chini et al., (2022) and Shah et al., (2024). Secondly, they show in the more strongly stratified case that these small horizontal scales are only dominant within the turbulent patches, and essentially disappear outside of these patches. As such, the distinct MSA model scalings are only expected to apply within the turbulent patches. In the more orderly layer-like flow outside of those patches, the SSA scalings—which coincide with the MSA model predictions in regions where small scales are not excited—should hold.
@@
To verify this interpretation quantitatively, we sought to identify a reliable diagnostic for the turbulent patches, i.e. regions where the flow exhibits small horizontal scales. It is common to use the enstrophy $|{\boldsymbol{\omega}}|^{2}$ as a diagnostic for turbulence, where $\boldsymbol{\omega}=\nabla\times{\bf u}$ is the flow vorticity. Indeed, the turbulent cascade to small scales implies that enstrophy must be large within the patches. However, enstrophy turns out to be an inappropriate diagnostic for our purpose because it can also be large in the layer-like regions of strong vertical shear outside of the turbulent patches, such as the ones described by the SSA model. This fact is illustrated in figure 4(a), which shows the enstrophy field in a particular snapshot of a strongly stratified simulation, and can be understood as follows. According to Chini et al., (2022) and Shah et al., (2024), ${\boldsymbol{u}}=\bar{\boldsymbol{u}}+{\boldsymbol{u}}^{\prime}$ where $\bar{\boldsymbol{u}}$ can be thought of as the large-scale anisotropic component of the flow, which varies on the $O(1)$ horizontal scales and $O(\alpha)$ vertical scale, as in the SSA model. Meanwhile ${\boldsymbol{u}}^{\prime}$ can be thought of as the small-scale isotropic and turbulent component of the flow, which varies on $O(\alpha)$ scales in all directions, as in the MSA model. Furthermore, these authors show that $\bar{u}\sim\bar{v}\sim O(1)$, while $\bar{w}\sim O(\alpha)$, and $u^{\prime}\sim v^{\prime}\sim w^{\prime}\sim O(\alpha^{1/2})$. Accordingly, we find that the horizontal vorticity components are dominated by the contribution from $\bar{\boldsymbol{u}}$, namely $\omega_{x}\sim\bar{\omega}_{x}\sim\omega_{y}\sim\bar{\omega}_{y}\sim O(Fr^{-1})$, while the vertical vorticity component is dominated by the contributions from ${\boldsymbol{u}}^{\prime}$, with $\omega_{z}\sim\omega^{\prime}_{z}\sim O(Fr^{-1/2})$. As a result, we argue that $\omega_{z}^{2}$ is a more reliable diagnostic of the small-scale turbulence than the enstrophy. This assertion is confirmed in figure 4(b), which shows $\omega_{z}^{2}$ for the same snapshot depicted in figure 4(a). We see that the regions of high $\omega_{z}^{2}$ only highlight the turbulent patches of the flow.
@@
In what follows, we therefore define the following quantities:
@@
The first can essentially be viewed as the root-mean-square of $w$ taken over the turbulent patches, where the distinct MSA scalings should apply. The second can be viewed as the rms of $w$ taken everywhere other than the turbulent patches, where the SSA scalings should apply. Note that the computation of $w^{\rm turb}_{rms}$ and $w^{\rm noturb}_{rms}$ requires integrals of $w^{2}$, $\omega_{z}^{2}$, and their product or ratio over the entire volume, which was not one of the simulation diagnostics originally saved. As such, we are unable to extract these quantities from the $Re=1000$ simulations. However, we can compute them from the full-data snapshots regularly saved in the $Re=600$ simulations in both high and low $Pe$ datasets (of which there are usually between 50 and 100 depending on the simulation). The variance is naturally larger than for the $w_{rms}$ data, owing to the intermittency of the turbulence.
@@ 4 Conclusion
In this paper, we have presented a detailed comparison of DNS data with various theoretical predictions for the characteristic vertical velocity of fluid motions in forced stratified turbulence. In particular, we have studied both moderate and low Prandtl number regimes, resulting in a wide range of Péclet numbers at fixed Reynolds number. When buoyancy diffusion is negligible, our results notably provide compelling evidence for the $w\propto Fr^{1/2}$ scaling law for stratified turbulence first proposed by Riley and Lindborg, (2012) using heuristic arguments and rigorously derived by Chini et al., (2022) using multiscale asymptotic analysis. In the latter investigation, this scaling law is intrinsically tied to the existence of small-scale isotropic flow motions driven by the emergent vertical shear between larger-scale primarily horizontal eddies, as corroborated by the results from the DNS presented here. The vertical shear instability is gradually stabilized as the buoyancy Reynolds number decreases towards a critical value of order unity, and the small-scale isotropic component of the turbulence becomes spatio-temporally intermittent rather than domain filling. Outside of the turbulent patches, small horizontal scales disappear, and we find that $w\propto Fr$ instead, consistent with the model of Billant and Chomaz, (2001) and Brethouwer et al., (2007). In this intermittent regime, therefore, the rms vertical velocity of the flow computed from an average over the whole domain differs from either of these scaling laws, and additionally depends on the volume filling factor of the small-scale turbulence, whose dependence on stratification will be the subject of future work. This intermittency also explains the incorrect conclusion reached by Garaud, (2020) regarding the possible existence of another regime of stratified turbulence where $w\propto Fr^{2/3}$. In hindsight, we understand her empirically-inferred scaling simply as a consequence of the decrease in the volume filled by turbulent patches with increasing stratification at fixed $Re$.
@@
At low $Pr$, Shah et al., (2024) revised the predictions of Brethouwer et al., (2007) and Chini et al., (2022) to account for the effects of buoyancy diffusion. They showed that the presence of small isotropic motions implies that $w\propto(Fr^{2}/Pe)^{1/6}$, consistent with an early model and DNS data by Cope et al., (2020), while in their absence $w\propto(Fr^{2}/Pe)^{1/4}$, consistent with predictions from Lignières, (2020) and Skoutnev, (2023). Revisiting the very low $Pr$ DNS of Cope et al., (2020) in this new light, we have confirmed both scaling laws within and outside of the turbulent patches, respectively. Finally, the models of Chini et al., (2022) and Shah et al., (2024) also predict where in parameter space viscous effects become important. We have confirmed these predictions, too, with our DNS data.
@@
In summary, this investigation demonstrates that the combination of rigorous multiscale analysis (Chini et al.,, 2022; Shah et al.,, 2024) with idealized DNS (Cope et al.,, 2020; Garaud,, 2020, and new simulations presented here) can be a powerful tool to identify and validate scaling laws for stratified turbulence across different regions of parameter space. In future work, we will incorporate the effects of rotation and magnetic fields, which must be taken into account for a more realistic description of stratified turbulence in geophysical and astrophysical settings.
@@ Acknowledgements
The authors gratefully acknowledge the Geophysical Fluid Dynamics Summer School (NSF 1829864), particularly the 2022 and 2023 programs. K.S. acknowledges funding from the James S. McDonnell Foundation. G.P.C. acknowledges funding from the U.S. Department of Energy through award DE-SC0024572. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.

### v2

@@ 2 Comparison of theory with DNS
To gain insight into the applicability of the predicted scalings, we examine the actual flow field more closely. The top three rows of figure 3 show snapshots of $u$ and $w$ in three different high-resolution DNS at $Pe=100$ and $Re=1000$, with $Fr$ decreasing from about $0.18$ to about $0.058$. It is clear that while the $Fr\simeq 0.18$ case is fully turbulent, the more strongly stratified $Fr\simeq 0.058$ case is not as the turbulence is localized to small ‘patches’ (see e.g. the regions of high $|w|$). Similar findings were reported by Cope et al. (2020) in their low $Pe$ simulations in the regime that they named ‘stratified intermittent’ (see their figure 6) and are also evident in Garaud (2020); see the volume renderings of $u$ and $w$ at $Fr=0.05$ ($B=400$) in her figure 1 for instance. The bottom row of figure 3 shows the kinetic energy spectra of the horizontal flows (red lines) and of the vertical flow (blue lines) for the same simulations. Different lines correspond to different instants in time, in order to illustrate the instrinsic variability of the spectra.
Figure 3: Top 3 rows: DNS snapshots of $u$ and $w$ in the $y=0$ plane for $Re=1000$, $Pe=100$ and various Froude numbers, with stratification increasing from top to bottom. Bottom row: Kinetic energy spectra of the horizontal flows (red lines) and of the vertical flows (blue lines) as a function of the horizontal wavenumber $k_{h}$, for the same three Froude numbers, with stratification increasing from left to right. Each line corresponds to a particular instant in time.
These results clearly illustrate the coexistence of large and small horizontal scales, with $u$ increasingly dominated by large scales with subdominant small scales as stratification increases, and $w$ increasingly dominated by small scales with subdominant large scales (cf. Riley & Lindborg, 2012). We see from the spectra in particular that the large scales are highly anisotropic. The horizontal flows have a kinetic energy spectrum proportional to $k_{h}^{-3}$ (where $k_{h}$ is the horizontal wavenumber), consistent with oceanic observations (e.g. Klymak & Moum, 2007; Falder et al., 2016) and the classical empirical ‘Garrett-Munk’ spectrum for internal waves (Garrett & Munk, 1975). It is also consistent with observations of turbulence in the stratosphere on $\mathit{O}$(10km) scales (Lilly & Lester, 1974). Meanwhile, the kinetic energy spectrum of vertical motions is a weakly increasing function of $k_{h}$ at large scales, which peaks at a value of $k_{h}$ that appears to scale as $Fr^{-1}$ (with the limited data available). This tentatively shows that $l_{z}^{*}\simeq FrL^{*}$ is indeed the injection scale for the vertical motions in these non-diffusive strongly-stratified simulations, consistent with the interpretation that they arise from shear instabilities of the layerwise horizontal flow on that scale.
At smaller scales, we see that the flow becomes much more isotropic. In the $Fr\simeq 0.18$ and $Fr=0.1$ cases, for which $Re_{b}=Fr^{2}Re\simeq 33$ and $10$, respectively, the isotropic small-scale flow seems to have a standard $k_{h}^{-5/3}$ spectrum until viscous effects come into play. In the $Fr\simeq 0.058$, by contrast, the spectrum of the small-scale flow is steeper than $k_{h}^{-5/3}$, indicating that the turbulence is somewhat suppressed. This observation is not surprising, because at these parameter values $Re_{b}\simeq 3.3$, and so viscous effects are likely to suppress inertial range dynamics that are unaffected by stratification, as well as suppressing the local vertical shear instability except in regions where the shear is exceptionally strong, leading to the patchiness of the flow observed in the snapshots.
@@
The snapshots in the more strongly stratified case further reveal that the small horizontal scales are only dominant within the turbulent patches and essentially disappear outside of these patches. As such, the distinct MSA model scalings are only expected to apply within the turbulent patches. In the more orderly layer-like flow outside of those patches, the SSA scalings—which coincide with the MSA model predictions in regions where small scales are not excited—should hold.
@@
To verify this interpretation quantitatively, we sought to identify a reliable diagnostic for the turbulent patches, i.e. regions where the flow exhibits small horizontal scales. It is common to use the enstrophy $|{\boldsymbol{\omega}}|^{2}$ as a diagnostic for turbulence, where $\boldsymbol{\omega}=\nabla\times{\bf u}$ is the flow vorticity. Indeed, the turbulent cascade to small scales implies that enstrophy must be large within the patches. However, enstrophy turns out to be an inappropriate diagnostic for our purpose because it can also be large in the layer-like regions of strong vertical shear outside of the turbulent patches, such as the ones described by the SSA model. This fact is illustrated in figure 4(a), which shows the enstrophy field in a particular snapshot of a strongly stratified simulation, and can be understood as follows. According to Chini et al. (2022) and Shah et al. (2024), ${\boldsymbol{u}}=\bar{\boldsymbol{u}}+{\boldsymbol{u}}^{\prime}$ where $\bar{\boldsymbol{u}}$ can be thought of as the large-scale anisotropic component of the flow, which varies on the $O(1)$ horizontal scales and $O(\alpha)$ vertical scale, as in the SSA model. Meanwhile ${\boldsymbol{u}}^{\prime}$ can be thought of as the small-scale isotropic and turbulent component of the flow, which varies on $O(\alpha)$ scales in all directions, as in the MSA model. Furthermore, these authors show that $\bar{u}\sim\bar{v}\sim O(1)$, while $\bar{w}\sim O(\alpha)$, and $u^{\prime}\sim v^{\prime}\sim w^{\prime}\sim O(\alpha^{1/2})$. Accordingly, we find that the horizontal vorticity components are dominated by the contribution from $\bar{\boldsymbol{u}}$, namely $\omega_{x}\sim\bar{\omega}_{x}\sim\omega_{y}\sim\bar{\omega}_{y}\sim O(Fr^{-1})$, while the vertical vorticity component is dominated by the contributions from ${\boldsymbol{u}}^{\prime}$, with $\omega_{z}\sim\omega^{\prime}_{z}\sim O(Fr^{-1/2})$. For comparison, note that the vertical vorticity associated with the large-scale forcing, and with the mean horizontal flow $(\bar{u},\bar{v})$, is $O(1)$ and therefore negligible in comparison with $\omega_{z}^{\prime}$. As a result, we argue that $\omega_{z}^{2}$ is a more reliable diagnostic of the small-scale turbulence than the enstrophy. This assertion is confirmed in figure 4(b), which shows $\omega_{z}^{2}$ for the same snapshot depicted in figure 4(a). We see that the regions of high $\omega_{z}^{2}$ only highlight the turbulent patches of the flow.
@@
In what follows, we therefore define the following quantities, using weighted averages with the weight function $\omega_{z}^{2}$ to emphasize the turbulent patches and the function $\omega_{z}^{-2}$ to emphasize the more quiescent regions, respectively:
@@
The first can be viewed as the rms of $w$ taken over the turbulent patches, where the distinct MSA scalings should apply. The second can be viewed as the rms of $w$ taken everywhere other than the turbulent patches, where the SSA scalings should apply. Note that the computation of $w^{\rm turb}_{rms}$ and $w^{\rm noturb}_{rms}$ requires integrals of $w^{2}$, $\omega_{z}^{2}$, and their product or ratio over the entire volume, which was not one of the simulation diagnostics originally saved. As such, we are unable to extract these quantities from the $Re=1000$ simulations. However, we can compute them from the full-data snapshots regularly saved in the $Re=600$ simulations in both high and low $Pe$ datasets (of which there are usually between 50 and 100 depending on the simulation). The variance is naturally larger than for the $w_{rms}$ data, because of the smaller amount of data available. We also note that similar results can be obtained by choosing $|\omega_{z}|$ and $|\omega_{z}|^{-1}$ as the weight functions for the turbulent and quiescent regions, respectively. However, $|\omega_{z}|$ is only a factor of $Fr^{-1/2}$ larger in the turbulent regions than in the quiescent ones. Since $Fr$ in our simulations is not extremely small, we use $\omega_{z}^{2}=O(Fr^{-1})$ instead to more clearly identify the turbulent patches.
@@ 4 Conclusion
In this paper, we have presented a detailed comparison of DNS data with various theoretical predictions for the characteristic vertical velocity of fluid motions in forced stratified turbulence. In particular, we have studied both moderate and low Prandtl number regimes, resulting in a wide range of Péclet numbers at fixed Reynolds number. When buoyancy diffusion is negligible, our results notably provide compelling evidence for the $w\propto Fr^{1/2}$ scaling law for stratified turbulence first proposed by Riley & Lindborg (2012) using heuristic arguments and rigorously derived by Chini et al. (2022) using multiscale asymptotic analysis. In the latter investigation, this scaling law is intrinsically tied to the existence of small-scale isotropic flow motions driven by the emergent vertical shear between larger-scale primarily horizontal eddies, as corroborated by the results from the DNS presented here. The vertical shear instability is gradually stabilized as the buoyancy Reynolds number decreases towards a critical value of order unity, and the small-scale isotropic component of the turbulence becomes confined to localized patches rather than being domain filling. Outside of these turbulent patches, small horizontal scales disappear, and we find that $w\propto Fr$ instead, consistent with the model of Billant & Chomaz (2001) and Brethouwer et al. (2007). As $Re_{b}$ gradually approaches unity from above, therefore, the rms vertical velocity of the flow computed from an average over the whole domain differs from either of these scaling laws, and additionally depends on the volume filling factor of the small-scale turbulence, whose dependence on stratification and $Re_{b}$ is the subject of ongoing work. The patchiness of the small-scale isotropic flow also explains the incorrect conclusion reached by Garaud (2020) regarding the possible existence of another regime of stratified turbulence where $w\propto Fr^{2/3}$. In hindsight, we understand her empirically-inferred intermediate scaling as a consequence of the decrease in the volume filled by turbulent patches with increasing stratification at fixed $Re$.
@@
At low $Pr$, Shah et al. (2024) revised the predictions of Brethouwer et al. (2007) and Chini et al. (2022) to account for the effects of buoyancy diffusion. They showed that the presence of small isotropic motions implies that $w\propto(Fr^{2}/Pe)^{1/6}$, consistent with an early model and DNS data by Cope et al. (2020), while in their absence $w\propto(Fr^{2}/Pe)^{1/4}$, consistent with predictions from Lignières (2020) and Skoutnev (2023). Revisiting the very low $Pr$ DNS of Cope et al. (2020) in this new light, we have confirmed both scaling laws within and outside of the turbulent patches, respectively. Finally, the models of Chini et al. (2022) and Shah et al. (2024) also predict where in parameter space viscous effects become important. We have confirmed these predictions, too, with our DNS data.
An important caveat of our conclusions is that the simulations were limited to a specific type of horizontal forcing. We believe that the results ought to apply more generally as long as the forcing drives primarily horizontal flows on long time scales and large length scales, but this conjecture will need to be verified in future work.
@@
In summary, this investigation demonstrates that the combination of rigorous multiscale analysis (Chini et al., 2022; Shah et al., 2024) with idealized DNS (Cope et al., 2020; Garaud, 2020, and new simulations presented here) can be a powerful tool to identify and validate scaling laws for stratified turbulence across different regions of parameter space. In future work, we will incorporate the effects of rotation and magnetic fields, which must be taken into account for a more realistic description of stratified turbulence in geophysical and astrophysical settings.
@@ Acknowledgements
This work uses the Expanse supercomputer at the San Diego Supercomputing Center. P. Garaud thanks the SDSC support team for their help. The authors gratefully acknowledge the Geophysical Fluid Dynamics Summer School (NSF 1829864), particularly the 2018, 2022 and 2023 programs. K.S. acknowledges funding from the James S. McDonnell Foundation. G.P.C. acknowledges funding from the U.S. Department of Energy through award DE-SC0024572. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.
@@
Appendix: Numerical considerations
The results presented in this paper require simulations that have achieved a statistically stationary state lasting at least 100 time units. This duration was chosen to ensure that the time series from which flow statistics are computed are sufficiently uncorrelated in time. Indeed, assuming a mean streamwise velocity of 1, the fluid has time to flow approximately 8 times through the domain in 100 time units when the domain length is $4\pi$. In practice, the mean streamwise velocity ranges from about 2 to 4, depending on the input parameters, and the integration time interval is often larger than 100 time units, so the true number of ‘laps’ is generally much higher. Time series of the instantaneous rms streamwise and vertical velocity for a few selected simulations in both non-diffusive and diffusive regimes are shown in figure 6.
Figure 6: Sample time series of the instantaneous rms horizontal velocity $u_{rms}$ (red lines) and vertical velocity $w_{rms}$ (blue lines), as a function of time for various simulations. Note that the starting points of the simulations have been offset to an arbitrary position for ease of visualization. The grey area shows the time-averaging interval used in figures 2 and 5.
We have also verified that the scaling laws obtained are independent of the selected domain size by performing a few simulations in a $8\pi\times 4\pi\times 2\pi$ domain with the same body force. With that choice, the domain is sufficiently wide to accommodate two wavelengths of the applied sinusoidal force in the spanwise direction, and two wavelengths of the fastest-growing mode of horizontal shear instability in the streamwise direction. Because of the heavily increased computational cost, we have only run cases for $Re=600,Pe=60$, and three values of $Fr$, and these simulations have been run for a shorter duration. Yet, as demonstrated in figure 7, the same scaling laws are found in the turbulent patches ($w^{\rm turb}_{rms}\propto Fr^{1/2}$) and outside of the turbulent patches ($w^{\rm noturb}_{rms}\propto Fr$), respectively. We note that the prefactor is slightly smaller, but this is not surprising given that the meandering flow structure is allowed to be a little different in the larger domain.
Figure 7: Left: Filled symbols show $w^{\rm turb}_{rms}$ and $w^{\rm noturb}_{rms}$ extracted from simulations in larger computational domains (of size $8\pi\times 4\pi\times 2\pi$) and can be compared to those extracted from simulations in regular-sized domains, shown as open symbols. Right: Snapshot of the $(x,y)$ plane at some arbitrary value of $z$ in the statistically stationary state of a simulation at $Re=600$, $Pe=60$, $Fr=0.1$, in a domain of size $8\pi\times 4\pi\times 2\pi$.